# FlatQuant: Flatness Matters for LLM Quantization

### v2

@@ Abstract
Recently, quantization has been widely used for the compression and acceleration of large language models (LLMs). Due to the outliers in LLMs, it is crucial to flatten weights and activations to minimize quantization error with the equally spaced quantization points. Prior research explores various pre-quantization transformations to suppress outliers, such as per-channel scaling and Hadamard transformation. However, we observe that these transformed weights and activations can still remain steep and outspread. In this paper, we propose FlatQuant (Fast and Learnable Affine Transformation), a new post-training quantization approach to enhance flatness of weights and activations. Our approach identifies optimal affine transformations tailored to each linear layer, calibrated in hours via a lightweight objective. To reduce runtime overhead, we apply Kronecker decomposition to the transformation matrices, and fuse all operations in FlatQuant into a single kernel. Extensive experiments show that FlatQuant sets up a new state-of-the-art quantization benchmark. For instance, it achieves less than 1% accuracy drop for W4A4 quantization on the LLaMA-3-70B model, surpassing SpinQuant by $\textbf{7.5}\%$. For inference latency, FlatQuant reduces the slowdown induced by pre-quantization transformation from 0.26x of QuaRot to merely 0.07x, bringing up to 2.3x speedup for prefill and 1.7x speedup for decoding, respectively. Code is available at: https://github.com/ruikangliu/FlatQuant.
@@ 1 Introduction
Nevertheless, it is non-trivial to get a flat distribution of weights and activations in LLMs. LLMs are known to have extreme outliers over activations (Dettmers et al., 2022; Xiao et al., 2023) and pivot tokens (Liu et al., 2024a; Sun et al., 2024). To alleviate this problem, various pre-quantization transformations are proposed to mitigate the impact of outliers (Xiao et al., 2023; Ashkboos et al., 2024; Liu et al., 2024b; Ma et al., 2024). However, we revisit these transformations and find them still sub-optimal in promoting flatness. For instance, per-channel scaling (Xiao et al., 2023; Shao et al., 2023) aims to balance the outliers between weights and activations, but it falls short of distributing outliers over the non-outlier channels. The recent Hadamard transformation (Ashkboos et al., 2024; Lin et al., 2024) attempts to solve this problem, while the individual characteristics of each linear layer are not considered. Moreover, the linear transformation introduced by these methods (Ma et al., 2024; Ashkboos et al., 2024; Liu et al., 2024b) inevitably introduces extra inference overhead that affects the overall speedup of quantization.
In this work, we introduce FlatQuant, a new post-training quantization approach. The name has dual significance: it stands for the proposed method (Fast and Learnable Affine Transformation) and emphasizes its goal of achieving flatter distributions of weights and activations, which are crucial for effective quantization. FlatQuant aims to identify the optimal affine transformation for each linear layer, employing a lightweight, block-wise training strategy over the calibration data. To minimize the inference overhead associated with affine transformations, FlatQuant harnesses the efficiency of Kronecker decomposition, reducing both the memory and computational demands. Our approach is compatible with various quantization techniques such as learnable clipping, and can be applied to various quantization settings, e.g., weight-only quantization or KV cache quantization. Additionally, as the affine transformations in FlatQuant are memory bound, we further fuse these transformations and quantization into a single kernel, thereby minimizing the global memory access and kernel lunch overhead. Lastly, extensive experiments are conducted to assess FlatQuant across various tasks, including language modeling and question answering, using LLaMA-2/3 models ranging from 7B to 70B parameters. The empirical results demonstrate that our proposed approach surpasses current state-of-the-art methods in terms of both accuracy and inference latency.
@@
  • We have designed an efficient kernel that fuses affine transformation and quantization, reducing the additional latency caused by transformation from a 0.26x slowdown with QuaRot to only 0.07x. This enhancement gives up to 2.3x speedup for prefill and 1.7x speedup for decoding compared to the FP16 baseline.
@@ 2.1 Preliminaries on LLM Quantization
where $s$ is the quantization step size, $\Pi(\cdot)$ is the projection function and $\Omega(b)=\{0,1,...,2^{b}-1\}$ is the set of $b$-bit integer points. For simplicity of notation, we denote $\mathcal{Q}(\cdot)$ as the general quantization function in the rest of this paper. Quantizing weights enables memory time savings from weight loading from high-bandwidth memory (HBM) into the compute cores, and quantizing activations further reduces the computation, benefiting both the prefill and decoding stages of LLM inference.
As recent works suggest (Xiao et al., 2023; Shao et al., 2023; Xi et al., 2023), LLMs exhibit persistent outliers in activations, posing significant challenges for quantization. Various works have been proposed to suppress these outliers to improve the quantized LLMs. Two most commonly used methods are per-channel scaling (Xiao et al., 2023; Lin et al., 2023; Wei et al., 2023) and Hadamard transformation or its variants (Xi et al., 2023; Ashkboos et al., 2024; Lin et al., 2024).
@@ 2.2 The Flatness for Quantization
We examine existing pre-quantization transformations with a focus on their potential for flatness, a critical factor for effective quantization. Intuitively, by removing outliers, these transformations are expected to produce flat weights and activations that are conducive to quantization. Additionally, the quantization error propagated through the network is also expected to be low and flat. However, our empirical results indicate that current methods are limited in achieving the desired flatness. This is in contrast to the proposed FlatQuant, as will be elaborated in Section 3. We leave how FlatQuant learns flatness in Appendix 8.
@@
Flat tensors are intuitively easier to quantize after removing outliers, a.k.a tensors with low kurtosis (Chmiel et al., 2020; Li et al., 2024). Figure 1 displays the distributions of both the original and transformed weights and activations, sorted by the channel magnitudes in descending order. The flat weights and activations with horizontal envelopes are usually preferred by quantization. Compared with the original distributions, pre-quantization transformations can yield flatter activations (e.g., Figure 1(b)1(d)) but still with their limitations. Per-channel scaling flattens activations at the cost of steeper weight envelops (e.g., Figure 1(a)1(c)). While Hadamard transformation produces generally better flatness for both activations and weights than per-channel scaling, it still sometimes generates unsatisfactory weights and activations distributions (e.g., Figure 1(a),  1(b)). In contrast, FlatQuant consistently flattens both weights and activations.
@@ 3.1 Fast and Learnable Affine Transformation
We begin with applying FlatQuant for a standard linear layer, and will discuss its integration with the Transformer architecture in Section 3.2. The overview of FlatQuant is presented in Figure 3. A primary objective of FlatQuant is to find the best affine transformation for each linear layer to quantize. Ideally, given $\mathbf{Y}=\mathbf{X}\mathbf{W}^{\top}$, one can identify the optimal invertible matrix $\mathbf{P}^{*}\in\mathbb{R}^{n\times n}$ by
$\mathbf{P}^{*}=\arg\min_{\mathbf{P}}\|\mathbf{Y}-\mathcal{Q}(\mathbf{X}\mathbf{P})\mathcal{Q}(\mathbf{P}^{-1}\mathbf{W}^{\top})\|_{F}^{2},$ (2)
@@
as studied in (Ma et al., 2024). The weights $\mathbf{P}^{-1}\mathbf{W}^{\top}$ can be pre-computed offline akin to (Ashkboos et al., 2024). However, unlike Hadamard matrices that can be reused for all layers, storing individual $\mathbf{P}$ matrices for different linear layers is computationally expensive. In the forward pass, this approach doubles the computational cost and memory access for matrix multiplication. Additionally, it nearly doubles the model storage requirements. Thus, another key aspect of FlatQuant is to identify a fast alternative for the pre-quantization transformation.
Kronecker Decomposition.
@@
We propose to decompose the original $\mathbf{P}\in\mathbb{R}^{n\times n}$ into $\mathbf{P}=\mathbf{P}_{1}\otimes\mathbf{P}_{2}$, where $\mathbf{P}_{1}\in\mathbb{R}^{n_{1}\times n_{1}},\mathbf{P}_{2}\in\mathbb{R}^{n_{2}\times n_{2}}$ are invertible matrices in smaller sizes, and $n=n_{1}n_{2}$. Recall the vectorization trick of the Kronecker product, i.e., $\textrm{vec}(\mathbf{V})(\mathbf{P}_{1}\otimes\mathbf{P}_{2})=\textrm{vec}(\mathbf{P}_{1}^{\top}\mathbf{V}\mathbf{P}_{2})$ for some $\mathbf{V}\in\mathbb{R}^{n_{1}\times n_{2}}$, the matrix multiplication in Equation 2 can be re-written as
$\mathcal{Q}(\mathbf{X}\mathbf{P})\mathcal{Q}(\mathbf{P}^{-1}\mathbf{W}^{\top})=\mathcal{Q}(\mathbf{P}_{1}^{\top}\times_{1}\tilde{\mathbf{X}}\times_{2}\mathbf{P}_{2})\ \mathcal{Q}(\mathbf{P}_{1}^{-1}\times_{1}\tilde{\mathbf{W}}\times_{2}(\mathbf{P}_{2}^{-1})^{\top})^{\top}\ ,$ (3)
@@
where $\tilde{\mathbf{X}}\in\mathbb{R}^{k\times n_{1}\times n_{2}}$ and $\tilde{\mathbf{W}}\in\mathbb{R}^{m\times n_{1}\times n_{2}}$ are reshaped from $\mathbf{X}$ and $\mathbf{W}$ accordingly, and ×$\times_{i}$ denotes the reduction over the $i$-th axis. Note that both weights and activations are converted back to matrix before multiplication. Such decomposition can save the memory up to $n/2$ times, given that $\frac{n^{2}}{n_{1}^{2}+n_{2}^{2}}\leq\frac{n^{2}}{2n_{1}n_{2}}=\frac{n}{2}$, with the equality holds when $n_{1}=n_{2}=\sqrt{n}$. Moreover, the computation saving is $\sqrt{n}/2$ times with the same optimal condition. In practice, we select $n_{1}^{*},n_{2}^{*}=\arg\min(n_{1}+n_{2})$, s.t. $n_{1}n_{2}=n$ and $n_{1}\leq n_{2}$. For instance, the optimal configuration is $(n_{1}^{*},n_{2}^{*})=(64,128)$ for $n=8192$. We find such a strategy gives the best speedup without compromising performance, as will be detailed in Figure 6. The affine transformations are therein pretty light-weight: they only account for 2.61% of the overall computational FLOPs and 0.11MB extra memory when integrated with LLaMA-2-7B. More details of inference overhead are in Appendix C.5.
@@
To further reduce the potential outlier after the above transformation, we combine learnable clipping thresholds $\alpha_{w},\alpha_{a}\in(0,1)$ on both weight and activation for each linear layer, together with the KV cache. While previous studies (Jacob et al., 2018; Frantar et al., 2022; Ashkboos et al., 2024) demonstrate that grid search is valid to find reasonable clipping thresholds, we observe that learning the clipping thresholds yields better results. These parameters are layer-specific and can be jointly optimized with the linear transformation matrices $\mathbf{P}$ and scaling vector $\textrm{diag}(\bm{c})$.A sigmoid function is applied to constrain $\alpha_{w}$ and $\alpha_{a}$ within $(0,1)$.
@@
We are now ready for the training objective of FlatQuant. We follow post-training quantization and sequentially minimize the mean squared error (MSE) by quantization over a small amount of calibration data (e.g., 128 randomly sampled sentences) for each Transformer block. The training objective for the $l$-th Transformer block is
@@
where $\mathcal{F}_{l}(\cdot)$ and ^$\hat{\mathcal{F}}_{l}(\cdot)$ denote the original and the quantized Transformer block, $\Theta=\{\mathbf{P},\bm{c},\alpha_{a},\alpha_{w}\}$ is abbreviated for all learnable parameters within the block. The transformation matrices within a Transformer block will be explained in Section 3.2. To compute the matrix inversion in Equation 3 efficiently and accurately, we adopt the singular value decomposition together with automatic mixed precision. More details can be found in Appendix B.1. Note that we also experiment with training multiple Transformer blocks together but find similar performance at higher training costs. Finally, we remark that the sequential training with Equation 4 can effectively produce flat weights and activations in Figure 1, and reduce the error propagation along the Transformer blocks in Figure 2.
@@ 3.2 Integration with the Transformer Architecture
As depicted in Figure 3, we now illustrate the integration of FlatQuant with a Transformer block based on an LLaMA-like architecture. Following the conventional practices, we employ low-bit matrix multiplications for all linear layers, while keeping layer normalization layers, pre-quantization transformations, RoPE embeddings, and attention scores in FP16.
@@
Recall that QuaRot (Ashkboos et al., 2024) and SpinQuant (Liu et al., 2024b) modify the LayerNorm to RMSNorm and merge orthogonal transformations into preceding layers for efficiency. Nonetheless, the residual connection of the “pre-norm” architecture would constrain all Transformer blocks to share the same transformation after RMSNorm. Instead, FlatQuant preserves the LayerNorm, and allows the use of fast and learnable affine transformations in Section 3.1 after LayerNorm for different layers, thereby enhancing the expressiveness.

### v3

@@ Abstract
Recently, quantization has been widely used for the compression and acceleration of large language models (LLMs). Due to the outliers in LLMs, it is crucial to flatten weights and activations to minimize quantization error with equally spaced quantization points. Prior research explores various pre-quantization transformations to suppress outliers, such as per-channel scaling and Hadamard transformation. However, we observe that these transformed weights and activations can still exhibit steep and dispersed distributions. In this paper, we propose FlatQuant (Fast and Learnable Affine Transformation), a new post-training quantization approach that enhances the flatness of weights and activations. Our approach identifies optimal affine transformations for each linear layer, calibrated in hours via a lightweight objective. To reduce runtime overhead of affine transformation, we apply Kronecker product with two lightweight matrices, and fuse all operations in FlatQuant into a single kernel. Extensive experiments demonstrate that FlatQuant establishes a new state-of-the-art benchmark for quantization. For example, it achieves less than 1% accuracy drop for W4A4 quantization on the LLaMA-3-70B model, surpassing SpinQuant by 7.5%. Additionally, it provides up to 2.3x prefill speedup and 1.7x decoding speedup compared to the FP16 model. Code is available at: https://github.com/ruikangliu/FlatQuant.
flatness, post-training-quantiation, affine transformation, large language models, INT4 quantization
@@ 1 Introduction
Nevertheless, it is non-trivial to get a flat distribution of weights and activations in LLMs. LLMs are known to have extreme outliers over activations (Dettmers et al., 2022; Xiao et al., 2023) and pivot tokens (Liu et al., 2024a; Sun et al., 2024). To alleviate this problem, various pre-quantization transformations are proposed to mitigate the impact of outliers (Xiao et al., 2023; Ashkboos et al., 2024; Liu et al., 2024c; Ma et al., 2024). However, we revisit these transformations and find them still sub-optimal in promoting flatness. For instance, per-channel scaling (Xiao et al., 2023; Shao et al., 2023) aims to balance the outliers between weights and activations, but it falls short of distributing outliers over the non-outlier channels. The recent Hadamard transformation (Ashkboos et al., 2024; Lin et al., 2024) attempts to solve this problem, while the individual characteristics of each linear layer are not considered. Moreover, the linear transformation introduced by these methods inevitably introduces extra inference overhead that affects the overall speedup of quantization.
In this work, we introduce FlatQuant, a new post-training quantization approach. The name has dual significance: it stands for the proposed method (Fast and Learnable Affine Transformation) and emphasizes its goal of achieving flatter distributions of weights and activations, which are crucial for effective quantization. FlatQuant aims to identify the optimal affine transformation for each linear layer, employing a lightweight, block-wise training strategy over the calibration data. To minimize the inference overhead associated with affine transformations, FlatQuant harnesses the efficiency of Kronecker product with two lightweight matrices, reducing both the memory and computational demands. Our approach is compatible with various quantization techniques such as learnable clipping, and can be applied to various quantization settings, e.g., weight-only quantization or KV cache quantization. Additionally, as the affine transformations in FlatQuant are memory bound, we further fuse them together with quantization into a single kernel, minimizing the global memory access and kernel lunch overhead. Lastly, extensive experiments over various tasks (e.g., language modeling and question answering) and LLM families (e.g., LLaMA and Qwen) are conducted to assess FlatQuant. Our empirical results demonstrate that FlatQuant establishes new state-of-the-art quantization results w.r.t. both accuracy and inference latency.
@@
  • We also design an efficient kernel that fuses both affine transformation and quantization, leading to 2.3x prefill speedup and 1.7x decoding speedup under W4A4 quantization when compared to the FP16 baseline.
@@ 2.1 Preliminaries on LLM Quantization
where $s$ is the quantization step size, $\Pi(\cdot)$ is the projection function and $\Omega(b)=\{0,1,...,2^{b}-1\}$ is the set of $b$-bit integer points. For simplicity of notation, we denote $\mathcal{Q}(\cdot)$ as the general quantization function in the rest of this paper.
As recent works suggest (Xiao et al., 2023; Shao et al., 2023; Xi et al., 2023), LLMs exhibit persistent outliers in activations, posing significant challenges for quantization. Various works are proposed to suppress outliers to improve the quantized LLMs. Two commonly used methods are per-channel scaling (Xiao et al., 2023; Lin et al., 2023; Wei et al., 2023) and Hadamard transformation or its variants (Xi et al., 2023; Ashkboos et al., 2024; Lin et al., 2024).
@@ 2.2 The Flatness for Quantization
Flat tensors are intuitively easier to quantize after removing outliers, a.k.a tensors with low kurtosis (Chmiel et al., 2020; Li et al., 2024). Figure 1 displays the distributions of both the original and transformed weights and activations, sorted by the channel magnitudes in descending order. The flat weights and activations with horizontal envelopes are usually preferred by quantization. Compared with the original distributions, pre-quantization transformations can yield flatter activations (e.g., Figure 1(b)1(d)) but still with their limitations. Per-channel scaling flattens activations at the cost of steeper weight envelops (e.g., Figure 1(a)1(c)). While Hadamard transformation produces better flatness for both activations and weights than per-channel scaling, it still sometimes generates unsatisfactory distributions (e.g., Figure 1(a),  1(b)). In contrast, FlatQuant, as will be elaborated in Section 3, consistently flattens both weights and activations.
@@ 3 Method
We now introduce FlatQuant with an overview in Figure 3. The proposed method has dual significance: it employs fast and learnable affine transformations, which also produce flat weights and activations for easy quantization.
@@ 3.1 Fast and Learnable Affine Transformation
We first introduce FlatQuant for a standard linear layer, and will discuss its integration with the Transformer architecture in Section 3.2. A primary objective of FlatQuant is to find the best affine transformation for each linear layer to quantize. Ideally, given $\mathbf{Y}=\mathbf{X}\mathbf{W}^{\top}$, one can identify the optimal invertible matrix $\mathbf{P}^{*}\in\mathbb{R}^{n\times n}$ by
$\mathbf{P}^{*}=\arg\min_{\mathbf{P}}\|\mathbf{Y}-\mathcal{Q}(\mathbf{X}\mathbf{P})\mathcal{Q}(\mathbf{P}^{-1}\mathbf{W}^{\top})\|_{F}^{2},\vspace{-1ex}$ (2)
@@
as studied in (Ma et al., 2024). The weights $\mathbf{P}^{-1}\mathbf{W}^{\top}$ can be pre-computed offline akin to (Ashkboos et al., 2024). However, unlike Hadamard matrices that can be reused for all layers, it is computationally expensive to maintain individual $\mathbf{P}$ matrices for different linear layers. In the forward pass, this approach doubles the computational cost and memory access for matrix multiplication. Additionally, it nearly doubles the model storage requirements.
Kronecker Product.
@@
The key of FlatQuant is to use Kronecker product of two lightweight matrices as an efficient substitution of the large affine transformation matrix $\mathbf{P}\in\mathbb{R}^{n\times n}$. Specifically, we propose to construct two lightweight matrices, i.e. $\mathbf{P}=\mathbf{P}_{1}\otimes\mathbf{P}_{2}$, where $\mathbf{P}_{1}\in\mathbb{R}^{n_{1}\times n_{1}},\mathbf{P}_{2}\in\mathbb{R}^{n_{2}\times n_{2}}$ are invertible matrices in smaller sizes, and $n=n_{1}n_{2}$. Recall the vectorization trick of the Kronecker product, i.e., $\textrm{vec}(\mathbf{V})(\mathbf{P}_{1}\otimes\mathbf{P}_{2})=\textrm{vec}(\mathbf{P}_{1}^{\top}\mathbf{V}\mathbf{P}_{2})$ for some $\mathbf{V}\in\mathbb{R}^{n_{1}\times n_{2}}$, the matrix multiplication in Equation 2 can be re-written as
$\mathcal{Q}$ $(\mathbf{X}\mathbf{P})\mathcal{Q}(\mathbf{P}^{-1}\mathbf{W}^{\top})=$ (3)
$\mathcal{Q}(\mathbf{P}_{1}^{\top}\times_{1}\tilde{\mathbf{X}}\times_{2}\mathbf{P}_{2})\times\mathcal{Q}(\mathbf{P}_{1}^{-1}\times_{1}\tilde{\mathbf{W}}\times_{2}(\mathbf{P}_{2}^{-1})^{\top})^{\top},\vspace{-1ex}$
@@
where $\tilde{\mathbf{X}}\in\mathbb{R}^{k\times n_{1}\times n_{2}}$ and $\tilde{\mathbf{W}}\in\mathbb{R}^{m\times n_{1}\times n_{2}}$ are reshaped from $\mathbf{X}$ and $\mathbf{W}$ accordingly, and ×$\times_{i}$ denotes the reduction over the $i$-th axis. Note that both weights and activations are converted back to matrix before multiplication. This design can save the memory up to $n/2$ times, given that $\frac{n^{2}}{n_{1}^{2}+n_{2}^{2}}\leq\frac{n^{2}}{2n_{1}n_{2}}=\frac{n}{2}$, with the equality holds when $n_{1}=n_{2}=\sqrt{n}$. Moreover, the computation saving is $\sqrt{n}/2$ times with the same optimal condition. In practice, we select $n_{1}^{*},n_{2}^{*}=\arg\min(n_{1}+n_{2})$, s.t. $n_{1}n_{2}=n$ and $n_{1}\leq n_{2}$. For instance, the optimal configuration is $(n_{1}^{*},n_{2}^{*})=(64,128)$ for $n=8192$. We find such a strategy gives the best speedup without compromising performance, as shown in Figure 6. The affine transformations are therein pretty light-weight, e.g., they only take 2.61% of the overall computational FLOPs and 3.41MB extra memory for LLaMA-2-7B. More details of inference overhead are in Appendix B.2.
@@
To further reduce the potential outlier after the above transformation, we combine learnable clipping thresholds $\alpha_{w},\alpha_{a}\in(0,1)$ after sigmoid functions on both weight and activation for each linear layer, together with the KV cache. While previous studies (Jacob et al., 2018; Frantar et al., 2022; Ashkboos et al., 2024) demonstrate that grid search is valid to find reasonable clipping thresholds, we observe that learning the clipping thresholds yields better results. These parameters are layer-specific and can be jointly optimized with the linear transformation matrices $\mathbf{P}$ and scaling vector $\textrm{diag}(\bm{c})$.
@@
We follow post-training quantization and sequentially minimize the mean squared error (MSE) by quantization over a small amount of calibration data (e.g., 128 randomly sampled sentences) for each Transformer block. The training objective for the $l$-th Transformer block is
@@
where $\mathcal{F}_{l}(\cdot)$ and ^$\hat{\mathcal{F}}_{l}(\cdot)$ denote the original and the quantized Transformer block, $\Theta=\{\mathbf{P},\bm{c},\alpha_{a},\alpha_{w}\}$ is abbreviated for all learnable parameters within the block. The transformation matrices within a Transformer block will be explained in Section 3.2. To compute the matrix inversion in Equation 3 efficiently and accurately, we adopt the singular value decomposition together with automatic mixed precision. See Appendix B.1 for details. We also tried with training multiple Transformer blocks together but found similar performance at higher training costs. Finally, the training progress with Equation 4 leads to flat weights and activations, as will be shown in Section 4.4.
@@ 3.2 Integration with the Transformer Architecture
We now illustrate the integration of FlatQuant with a Transformer block based on an LLaMA-like architecture. Following the conventional practices, we employ low-bit matrix multiplications for all linear layers, while keeping layer normalization layers, pre-quantization transformations, RoPE embeddings, and attention scores in FP16.
@@
Recall that QuaRot (Ashkboos et al., 2024) and SpinQuant (Liu et al., 2024c) modify the LayerNorm to RMSNorm and merge orthogonal transformations into preceding layers for efficiency. Nonetheless, the residual connection of the “pre-norm” architecture would constrain all Transformer blocks to share the same transformation after RMSNorm. Instead, FlatQuant preserves the LayerNorm, and allows the use of fast and learnable affine transformations in Section 3.1 after LayerNorm for different layers, thereby enhancing the expressiveness.